Can someone help me by answering #2 of these questions.Using the techniques described in sections 10.2 and 10.3 of the text, for each of these hypothesis testing problems, (a) stat
Answer #1

2)

New Old Difference(d)
13.3 14 -0.7
14.9 10.1 4.8
13.9 8.7 5.2
6.4 17.9 -11.5
11.1 10.7 0.4
7.3 12.8 -5.5
8.9 15 -6.1
12.9 15 -2.1
7.9 18.3 -10.4
11.5 15.4 -3.9
Total -29.8

a)

Claim: The new machine faster than the old machine.

The null and alternative hypothesis is

H0:\mu_{d}=0

H1:\mu_{d}>0

b)

Level of significance = 0.05

Degrees of freedom = n - 1 = 10 - 1 = 9

Critical value = 1.833   ( Using t table)

c)

Sample size = n = 10

Sample mean of difference = \bar{d} = - 2.98

Sample standard deviation of difference = s_{d} = 5.6707

Test statistic is

t=\frac{\bar{d}}{s_{d}/\sqrt{n}}

一2.98 5.6707/YW =-1.66
d)

Test statistic | t | < critical value we fail to reject the null hypothesis.

e)

Conclusion:

The new machine NOT faster than the old machine.